the distance from brampton to peterborough is 240 km. Chris is able to drive at an average speed of 100km/h along highway 401 and 80km/h on highway 115. If the total trip takes 2.5h, how long does he drive alog highway 115?
Found 2 solutions by rwm, htmentor:
Answer by rwm(914)

100*x+80*y=240
x+y=2.5
x=2 y=.5
half an hour on rte 115

Answer by htmentor(809)

the distance from brampton to peterborough is 240 km. Chris is able to drive at an average speed of 100km/h along highway 401 and 80km/h on highway 115. If the total trip takes 2.5h, how long does he drive alog highway 115?
===================================================================
Let t = the time traveled on highway 115
Since the total time is 2.5 h, the time traveled on highway 401 will be 2.5 - t
Since distance = speed x time, we can write
240 km = 80 km/h *t + 100 km/h *(2.5-t)
Simplify and solve for t:
240 = 80t + 250 - 100t
20t = 10
t = 0.5 h
So the time traveled along highway 115 is 0.5 hours
Check:
80*0.5 + 100*(2.5-0.5) = 40 + 200 = 240